Topic: Simultaneuos equations
If possible i would like a step by step guide of the answer thanks.
x+y=9
x^2-3xy+2y^2=0

Answer by checkley71(8403)

FIRST YOU NEED TO DEFINE X=9-Y THEN REPLACE X BY (9-Y) IN THE OTHER EQUATION
X^2-3XY+2Y^2=0
(X-2Y)(X-Y)=0
9-Y-2Y=0
9-3Y=0
-3Y=-9
Y=-9/-3
Y=3 ANSWER.
THEN X=9-3=6 ANSWER.
9-Y-Y=0
9-2Y=0
-2Y=-9
Y=-9/-2
Y=4.5 ANSWER.
THEN X=9-4.5=4.5 ANSWER.